# The illusion of households as entities in social networks

### v1

@@ 1 Introduction
Plenty of social science research ranging from within sociology (Campbell, 1958), political science (Nickerson, 2008; Auld and Zhang, 2013; Cheema et al., 2023), anthropology (Werner, 1998; Schmink, 1984; Niehof, 2011), and network science (Kumar et al., 2024) have theorized, measured, and observed that there are assumptions with consequences when studying individuals and their relationships interchangeably with their respective aggregates.
@@ 2 Related Work
The study of how individuals within social networks act and react in aggregate broadly spans political science, anthropology, and sociology. We briefly introduce this related work so that we may adapt it to make concrete recommendations the context of social networks in Section 5. We note that the related work we consider here often spans or defies disciplinary boundaries. As considered in the network science literature, the abstract question we consider (of node-aggregation) is closely related to the literature on coarse-graining (Itzkovitz et al., 2005; Gfeller and De Los Rios, 2007; Kim, 2004; Klein and Hoel, 2020) where the goal is to explore how to aggregate nodes while preserving some property of the network (e.g., degree distribution, spectral properties, properties of random walks). However, as we will see in this section, questions of how people and their relationships can be considered in aggregates have been considered in many disparate literatures.
@@
Household strategies
@@ 3.1 Node contraction
Household networks are typically studied by aggregating information from the individual network level using the basic households rule in Equation 1 (see LABEL:tab:review). However, the household strategies literature we briefly reviewed in Section 2 highlights that there is a key difference between individual relationships and those which entire households utilize, benefit from, and maintain. The contraction rules previously discussed are most suitable in contexts wherein it is appropriate to construct the household network from the individual network—as opposed to constructing the household network from particular household generating questions. This distinction is highlighted by contrasting the implications of two common name generator questions, “Who would you borrow kerosene or rice from?” and “Who do you go to for medical advice?” (e.g., Banerjee et al., 2013). The first question may generate names of individuals who are stand-ins for an entire household (if you go to X’s house to ask for rice but X is not home, do you ask whoever answers the door?). The latter question, however, generates a particular individual network, distinct from a shared household one (the trust involved in asking X for advice about your health may not transfer to X’s spouse).
@@ 4.1 Random graphs with random households
To understand how node contraction can impact the structure of the resulting graph, we first explore an example of randomly contracting nodes in a random graph on $n$ nodes where an edge exists between any two nodes with independent probability $p$, called an Erdős-Rényi random graph and denoted $G(n,p)$. In Appendix A we show that when households are of the exact same size, the resulting graph can be explained by the same Erdős-Rényi generative process, but fewer nodes and with a different edge probability. This observation, however, no longer holds when we consider households of differing sizes.
@@ 4.2 Local network properties
Figure 4: The inversity of the Banerjee et al. (2013) networks at the individual (blue) and household (red) level. In grey, we plot the histogram of the inversity of the individual networks across villages when the intrahousehold edges are removed.
@@ 4.3 Spread of information over a network
for $p\in[0,1]$. Here, $p$ can allow for flexibility in the presence of relationships between people in the same household, by varying the probability that an intrahousehold connection exists in a given context. For example, small $p$ represents a high probability that information, once given to one household member, will spread to the other members of the household. By contrast, large $p$ would capture a situation where there is low probability of information spreading between household members. By making this distinction, we build upon the modeling recommendations of Newman and Park (2003), wherein they consider how to project bipartite affiliation networks onto a corresponding social network. In the context of modeling the spread of information, writing adjacency matrices in this way can represent a rate of the spread of information within households and make a direct modeling connection to the work of, e.g., Nickerson (2008) who found that information does not deterministically reach all household members.
@@ 4.4 Diffusion centrality and gendered networks
Diffusion centrality, a variation of eigenvector centrality, can be interpreted as how extensively a piece of information spreads as a function of an initially informed node. Consider a directed, strongly connected network with corresponding weighted adjacency matrix $\boldsymbol{w}$ where $w_{ij}$ gives the relative probability that $i$ shares information with $j$. Then for a piece of information originating at node $i$, in each time period an informed node tells each neighbour $j$ the information with independent probability $w_{ij}$. If we consider that this process happens for $0<T<\infty$ time periods, diffusion centrality is specifically defined as:
@@ 5.1 Entitativity criteria
Proximity of individuals is determined by their physical closeness to one another. Although it may seem straightforward to assess this criteria for individuals in a household (because they all live in the same house), in some places and on certain timescales, household members may actually not be proximate. For example, assessing proximity must be approached with care in places where it’s common for one or more household members to leave the village/town for work for extended periods of time (see Niehof, 2011), or where the boundary of the household is more culturally nebulous (Yotebieng and Forcone, 2018). Assessing this question in an experimental context depends both on the timescale of the intervention and on cultural norms regarding household membership. These considerations aside, an aggregate of individuals living in the same household necessarily meet this criterion. Assessing if the household meets the following entitativity criteria alongside proximity considers whether or not the household boundary is what Campbell calls an illusion—a boundary that satisfies one criterion but not others.
@@
Similarity of individuals within their household unit should be particularly assessed with respect to an experimental question and/or intervention. Although not exhaustive, helpful dimensions along which to assess similarity are gender, power, and household roles. For example, if the experiment is focused on the spread of information, similarity must be assessed with respect to the content of the information that the individuals are exposed to and if the individuals’ differences in gender, power, or household roles impact the sharing of information. Guiding questions to assess this criterion in a given context might be:
@@
  • Does the subject or outcome of the experiment lie primarily within the set of roles of one household member?
@@
We add this criterion to Campbell’s set of criteria for use in the particular setting of analyzing networks, and is helpful in cases where determining entitativity might be non-obvious. This criterion is meant to assess whether or not the particular network metric of interest (e.g., an influence-maximizing seed set, clustering coefficient, inversity, or diffusion centrality) is qualitatively consistent regardless of which network is used. If it is not, we recommend analysis of the individual network.
@@ 6 Conclusions
In our discussion of the entitativity criterion of proximity in Section 5.1, we mentioned the additional difficulty in assessing this criterion in the case of longitudinal network studies or in cultures where the household unit is more ambiguously defined. Rigorous recommendations for how to define the boundaries of a household when its members do not remain consistent over time is an avenue for practical and impactful future work. Furthermore, as mentioned earlier, we did not consider Campbell’s criterion of assessing a household’s resistance to intrusion. In many respects, we can consider the very act of network data collection as an act of intrusion. Scientists have long considered impacts of the observer effect: how the collection of data from a system impacts the system itself. With respect to social data, anthropologists center the consideration of how the observer effect can either invalidate, bias, or strengthen ethnographic research findings (LeCompte and Goetz, 1982; Monahan and Fisher, 2010). Considering how the intrusion of social network data collection impacts the communities being researched—and the subsequent data—and incorporating households’ reactions to that intrusion in evaluating entitativity, is an interesting direction for future work.
We hope that by detailing the choices, consequences, and recommendations for studying individual and household networks, this work serves to inspire a rigorous chain of questions about the data decisions that we make when we aim to use networks as tools for exploring experimental questions about communities and the social ties between the people within them.

### v2

@@ 1 Introduction
Plenty of social science research ranging from within sociology (Campbell, 1958), political science (Nickerson, 2008; Auld and Zhang, 2013; Cheema et al., 2023), anthropology (Werner, 1998; Schmink, 1984; Niehof, 2011; Koster, 2018), and network science (Kumar et al., 2024) have theorized, measured, and observed that there are assumptions with consequences when studying individuals and their relationships interchangeably with their respective aggregates.
@@ 2 Related Work
The study of how individuals within social networks act and react in aggregate broadly spans political science, anthropology, and sociology. We briefly introduce this related work so that we may adapt it to make concrete recommendations the context of social networks in Section 5. We note that the related work we consider here often spans or defies disciplinary boundaries. As considered in the network science literature, the abstract question we consider (of node-aggregation) is closely related to the literature on coarse-graining (Itzkovitz et al., 2005; Gfeller and De Los Rios, 2007; Kim, 2004; Klein and Hoel, 2020) where the goal is to explore how to aggregate nodes while preserving some property of the network (e.g., degree distribution, spectral properties, properties of random walks). However, as we will see in this section, questions of how people and their social interactions can be considered in aggregates have been considered in many disparate literatures.
@@
Epidemiology
Within epidemiology there has been a focus on developing variations of epidemiological models to account for differences in transmission rates between individuals living within the same household, and how accounting for households and higher-order interactions changes the predicted spread of a disease (Becker and Dietz, 1995; Fraser, 2007; St-Onge et al., 2021; Boccaletti et al., 2023). This literature contributes not only specific modeling choices to account for the impact of households on disease transmission, but also makes more general graph-theoretical contributions. In Section 4.3 we draw upon the modeling distinctions from this body of work in how we model the spread of information between individuals within and beyond those living in the same household. In contrast to this work, we consider how household networks are constructed, the implications of treating households as social entities in social (non-biological) processes, and under what conditions it is appropriate to do so.
@@
Household strategies and household decision making
@@
These ethnographic and theoretical findings are echoed in economic theory, where work shows that there is an economic advantage for a division of labor amongst members in a household (Becker, 1993; Lundberg, 2008). This economic literature on household decision-making departs from a more traditional economic model of the household as a single unit. By accounting for differences in the individuals within a household, they show that it is more efficient for the utility of a household as a whole if individuals specialize their attention and labor in either market or household decisions. Whereas the divide in decision-making responsibilities may be determined by gender, when there is any difference in the aptitudes, incentives, or interests of individuals within a household, specialization increases the efficiency of the household as a whole. These findings provide theoretical grounding that decisions, even if made at the household level, may be the primary responsibility of only one member of the household.
@@ 3.1 Node contraction
Household networks are typically studied by aggregating information from the individual network level using the basic households rule in Equation 1 (see LABEL:tab:review). However, the household strategies literature we briefly reviewed in Section 2 highlights that there is a key difference between individual relationships and those which entire households utilize, benefit from, and maintain. The contraction rules previously discussed are most suitable in contexts wherein it is appropriate to construct the household network from the individual network—as opposed to constructing the household network from particular household generating questions.
This distinction is highlighted by contrasting the implications of two common name generator questions, “Who would you borrow kerosene or rice from?” and “Who do you go to for medical advice?” (e.g., Banerjee et al., 2013). The first question may generate names of individuals who are stand-ins for an entire household (if you go to X’s house to ask for rice but X is not home, do you ask whoever answers the door?). The latter question, however, generates a particular individual network, distinct from a shared household one (the trust involved in asking X for advice about your health may not transfer to X’s spouse). Furthermore, how respondents interpret and answer these survey questions is further complicated by the ambiguity as to whether “you” is implied to signify the individual survey respondent or the household they represent. That is, in addition to questions soliciting nominees which represent an entire household, in some circumstances the respondent may be answering on behalf of their entire household. This ambiguity is dependent on the culture, language, and setting in which the question is being asked.
@@
That is, if one can contextually distinguish certain name generators as soliciting relationships between households, a household network can be reasonably constructed by only connecting households which are connected in those particular layers.
@@ 4.1 Random graphs with random households
To understand how node contraction can impact the structure of the resulting graph, we first explore an example of considering the effect of randomly contracting nodes in a random graph, formalizing a result regarding expected infectious contacts between households noted in Doenges et al. (2024). Consider a graph on $n$ nodes where an edge exists between any two nodes with independent probability $p$, called an Erdős-Rényi random graph and denoted $G(n,p)$. In Appendix A we show that when households are of the exact same size, the resulting graph can be explained by a Erdős-Rényi generative process, but with fewer nodes and with a different edge probability. This observation, however, no longer holds when we consider households of differing sizes.
@@ 4.2 Local network properties
Figure 4: The inversity of the Banerjee et al. (2013) networks at the individual and household level. In light grey, we plot the histogram of the inversity of the individual networks across villages when the intrahousehold edges are removed.
@@ 4.3 Spread of information over a network
for $p\in[0,1]$. Here, $p$ can allow for flexibility in the presence of relationships between people in the same household, by varying the probability that an intrahousehold connection exists in a given context. For example, small $p$ represents a high probability that information, once given to one household member, will spread to the other members of the household. By contrast, large $p$ would capture a situation where there is low probability of information spreading between household members.
By making this distinction, we build upon the modeling recommendations of Newman and Park (2003), wherein they consider how to project bipartite affiliation networks onto a corresponding social network. In the context of modeling the spread of information, writing adjacency matrices in this way can represent a rate of the spread of information within households and make a direct modeling connection to the work of, e.g., Nickerson (2008) who found that information does not deterministically reach all household members. Furthermore, our modeling choice to allow for variation between the probability of information spreading between individuals within and beyond households is reflected in the epidemiological work considering how household transmission affects the spread of a disease (e.g., Ball and Sirl, 2012).
@@ 4.4 Diffusion centrality and gendered networks
Diffusion centrality, a variation of eigenvector centrality, as defined for directed networks in Banerjee et al. (2019), can be interpreted as how extensively a piece of information spreads as a function of an initially informed node. Consider a directed, strongly connected network with corresponding weighted adjacency matrix $\boldsymbol{w}$ where $w_{ij}$ gives the relative probability that $i$ shares information with $j$. Then for a piece of information originating at node $i$, in each time period an informed node tells each neighbour $j$ the information with independent probability $w_{ij}$. If we consider that this process happens for $0<T<\infty$ time periods, diffusion centrality is specifically defined as:
@@ 5.1 Entitativity criteria
Proximity of individuals is determined by their physical closeness to one another. Although it may seem straightforward to assess this criteria for individuals in a household (because they all live in the same house), in some places and on certain timescales, household members may actually not be proximate. For example, assessing proximity must be approached with care in places where it’s common for one or more household members to leave the village/town for work for extended periods of time (see Niehof, 2011), where the boundary of the household is more culturally nebulous (Yotebieng and Forcone, 2018), or when household and family composition changes in longitudinal studies (Steele et al., 2019). Determining the boundaries of what defines a household, who is in a household, and whether household members are proximate in an experimental context depends both on the timescale of the intervention and on cultural norms regarding household membership. These considerations aside, an aggregate of individuals living in the same household necessarily meet this criterion. Assessing if the household meets the following entitativity criteria alongside proximity considers whether or not the household boundary is what Campbell calls an illusion—a boundary that satisfies one criterion but not others.
@@
Similarity of individuals within their household unit should be particularly assessed with respect to an experimental question and/or intervention. Similarity here not only refers to a relevancy and accessibility of the intervention with respect to different household members’ identities (Berti et al., 2015), but also to their roles and responsibilities within the household (Becker, 1993). For example, if the experiment is focused on the spread of information, similarity must be assessed with respect to the content of the information that the individuals are exposed to and if the individuals’ differences in gender, power, or household responsibilities impact the sharing of information. Guiding questions to assess this criterion in a given context might be:
@@
  • Does the subject or outcome of the experiment lie primarily within the specialization of one household member?
@@
We add this criterion to Campbell’s set of criteria for use in the particular setting of analyzing networks, and is helpful in cases where determining entitativity might be non-obvious. This criterion is meant to assess whether or not the particular network metric of interest (e.g., an influence-maximizing seed set, clustering coefficient, inversity, or diffusion centrality) is qualitatively consistent regardless of which network is used. As an alternative to testing the consistency of a network metric, methodological work by Koster (2018) proposes a generalized linear mixed model for distinguishing household and individual effects in social networks. Employing this model as an evaluation of this criteria to statistically evaluate the extent to which individual intra-household relationships are shared between household members may indicate when aggregating connections within households is a representative assumption. In the case of inconsistent metrics, we recommend analysis of the individual network.
@@ 6 Conclusions
In our discussion of the entitativity criterion of proximity in Section 5.1, we mentioned the additional difficulty in assessing this criterion in the case of longitudinal network studies or in cultures where the household unit is more ambiguously defined. Rigorous recommendations for how to define the boundaries of a household when its members do not remain consistent over time has been meaningfully addressed in, e.g., Steele et al. (2019). Determining how to define a household network in a similar setting of changing household memberships and boundaries is an avenue for practical and impactful future work. Furthermore, as mentioned earlier, we did not consider Campbell’s criterion of assessing a household’s resistance to intrusion. In many respects, we can consider the very act of network data collection as an act of intrusion. Scientists have long considered impacts of the observer effect: how the collection of data from a system impacts the system itself. With respect to social data, anthropologists center the consideration of how the observer effect can either invalidate, bias, or strengthen ethnographic research findings (LeCompte and Goetz, 1982; Monahan and Fisher, 2010). Considering how the intrusion of social network data collection impacts the communities being researched—and the subsequent data—and incorporating households’ reactions to that intrusion in evaluating entitativity, is an interesting direction for future work.
We hope that by detailing the choices, consequences, and recommendations for studying individual and household networks, this work serves to encourage a rigorous chain of questions about the data decisions that we make when we aim to use networks as tools for exploring experimental questions about communities and the social ties between the people within them.